# Stable Gradients for Stable Learning at Scale in Deep Reinforcement Learning

### v1

@@ Abstract
Scaling deep reinforcement learning networks is challenging and often results in degraded performance, yet the root causes of this failure mode remain poorly understood. Several recent works have proposed mechanisms to address this, but they are often complex and fail to highlight the causes underlying this difficulty. In this work, we conduct a series of empirical analyses which suggest that the combination of non-stationarity with gradient pathologies, due to suboptimal architectural choices, underlie the challenges of scale. We propose a series of direct interventions that stabilize gradient flow, enabling robust performance across a range of network depths and widths. Our interventions are simple to implement and compatible with well-established algorithms, and result in an effective mechanism that enables strong performance even at large scales. We validate our findings on a variety of agents and suites of environments. We make our code publicly available.
@@ 1 Introduction
Recent advances in deep reinforcement learning (deep RL) have demonstrated the ability of deep neural networks to solve complex decision-making tasks from robotics to game play and resource optimization (Mnih et al., 2015, Vinyals et al., 2019, Bellemare et al., 2020, Fawzi et al., 2022). Motivated by successes in supervised and generative learning, recent works have explored scaling architectures in deep RL, showing gains in representation quality and generalization across tasks (Farebrother et al., 2023, Taiga et al., 2023). However, scaling neural networks in deep RL remains fundamentally challenging (Ceron et al., 2024b, a).
A central cause of this instability lies in the unique optimization challenges of RL. Unlike supervised learning, where data distributions are fixed, deep RL involves policy-dependent data that constantly change during training (Lyle et al., 2022). Each update of the policy $\pi_{\theta}$ alters future states and rewards, making the training objective inherently non-stationary. Value-based methods exacerbate these issues via bootstrapping, recursively using predicted values as targets. Estimation errors compound over time (Fujimoto et al., 2018), especially under sparse or delayed rewards (Zheng et al., 2018), leading to unstable updates, policy collapse, or value divergence (Van Hasselt et al., 2016, Lyle et al., 2023, 2024). These dynamics are tightly coupled with architectural vulnerabilities. Deep networks face well known pathologies such as vanishing/exploding gradients (Pascanu et al., 2013), ill-conditioned Jacobians (Pennington et al., 2017), and activation saturation (Glorot and Bengio, 2010). In deep RL, these are magnified by the “deadly triad(Sutton and Barto, 2018, Van Hasselt et al., 2018), off-policy corrections, and changing targets. As networks scale, the risk of signal distortion and misalignment increases, resulting in underutilized capacity and brittle learning (Obando Ceron et al., 2023, Ceron et al., 2024a).
One overlooked source of these failures lies in how gradients propagate through the network. Specifically, the gradient decomposition, the layer-wise structure of backpropagation as a chain of Jacobians and weights determine how information flows during learning (Lee et al., 2020). While gradient signal preservation has been studied in supervised learning (Schoenholz et al., 2017, Jacot et al., 2018), its role in deep RL, where both inputs and targets shift continually, remains poorly understood.
@@ 2 Preliminaries
A deep reinforcement learning agent interacts with an environment through sequences of actions ($a\in\mathcal{A}$), which produce corresponding sequences of observations ($s\in\mathcal{S}$) and rewards ($r\in\mathbb{R}$), resulting in trajectories of the form $\tau:=\{s_{0},a_{0},r_{0},s_{1},a_{1},r_{2},\ldots\}$. The agent’s behavior is often represented by a neural network with parameters $\theta$, composed of convolutional layers $\{\phi_{1},\phi_{2},\dots,\phi_{L_{c}}\}$ and dense (fully connected) layers $\{\psi_{1},\psi_{2},\dots,\psi_{L_{d}}\}$, where $\psi_{L_{d}}$ has an output dimensionality of $|\mathcal{A}|$. At every timestep $t$, an observation $s_{t}\in\mathcal{S}$ is fed through the network to obtain an estimate of the long-term value of each action: $Q_{\theta}(s_{t},\cdot)=\psi_{L_{d}}(\psi_{L_{d}-1}(\ldots(\phi_{L_{c}}(\ldots(\phi_{1}(s_{t}))\ldots))\ldots))$. The agent’s policy $\pi_{\theta}(\cdot\mid s_{t})$ specifies the probability of selecting each action, for instance by taking the softmax over the estimated values as in Equation 1. The training objective is typically defined as the maximization of expected cumulative reward as in Equation 2,
@@
While training large models in supervised learning settings present challenges, advances in initialization, normalization, and scaling strategies have enabled relatively stable optimization (Ioffe and Szegedy, 2015, Ba et al., 2016, Glorot and Bengio, 2010). Scaling up model size has been a central driver of progress across domains, improving generalization, enhancing representation learning, and boosting downstream performance (Kaplan et al., 2020).
Deep RL differs substantially from supervised learning. First, the data distribution is non-stationary, continually shifting as $\pi_{\theta}$ updates. Second, learning signals are often sparse, delayed, or noisy, which introduces variance in the estimated gradients (Han et al., 2022, Fujimoto et al., 2018). These factors destabilize optimization and lead to loss surfaces with sharp curvature and complex local structure (Ilyas et al., 2020, Achiam et al., 2019). Moreover, increasing model capacity often degrades performance unless regularization or architectural interventions are applied (Gogianu et al., 2021, Bjorck et al., 2021, Schwarzer et al., 2023, Wang et al., 2025).
@@
where each $\frac{\partial h_{k}}{\partial h_{k-1}}\in\mathbb{R}^{d_{k}\times d_{k-1}}$ is the Jacobian. If the singular values of these Jacobians are not properly controlled, their repeated multiplication can cause the norm of the gradient to shrink or grow exponentially with $\mathcal{L}$. This severely impairs convergence, as earlier layers receive little to no useful gradient signal or become numerically unstable (Ioffe and Szegedy, 2015, He et al., 2016).
@@
where $\zeta^{\prime}(\cdot)$ denotes the elementwise derivative of the activation function, and $\odot$ represents elementwise multiplication. For ReLU, $\zeta^{\prime}(x)=\mathbf{1}_{x>0}$, so the gradient is entirely blocked wherever the neuron is inactive. This leads to the well-known dying ReLU problem, where a significant portion of the network ceases to update and becomes untrainable (Lu et al., 2019, Shin and Karniadakis, 2020).
@@ 3 Diagnosis: Gradients Under Non-stationarity and Scale
A fundamental premise of modern deep learning is that scaling model capacity yields consistent gains in performance (Kaplan et al., 2020, Chowdhery et al., 2023). This has held true in large-scale supervised learning, where training data distributions are stationary and i.i.d., and gradient descent operates under relatively stable conditions. However, in non-stationary settings, such as RL, gradient-based optimization faces severe challenges that scaling alone may exacerbate (Ceron et al., 2024a, b). In this section, we diagnose how gradient pathologies emerge and intensify across different settings, with a focus on architectural scaling in width and depth (network scales used specified in Table 1).
@@ 3.1 Gradient Pathologies
Supervised Learning (Stationary and Non-Stationary) We use the CIFAR-10 image classification benchmark (Krizhevsky et al., 2009), where the input-output mapping remains fixed over time. Models consist of standard 6-layer convolutional neural networks (CNN) followed by a multi-layer perceptron (MLP). We vary the depth and width of the MLP to explore how model scale influences learning behavior. To introduce non-stationarity, we periodically shuffle the training labels during training, following the setup by Sokar et al. (2023). This creates a loss landscape that changes over time, echoing the challenges of deep RL. Figure 1 illustrates the contrast in training behavior and gradient flow between stationary and non-stationary supervised learning. Under non-stationarity, deep networks fail to recover accuracy, which aligns with a marked degradation in gradient magnitudes.
@@
Reinforcement Learning As discussed in section 2, RL introduces fundamentally different sources of non-stationarity due to the policy-dependent data distribution and moving target estimates. To study gradient dynamics, we use PQN (Gallici et al., 2025), a recent value-based algorithm that achieves strong performance without relying on a target network or replay buffer. PQN ensures stability and convergence using Layer Normalization (Ba et al., 2016) and supports GPU-based training through vectorized environments for online parallel data collection. In subsection C.1 we extend our investigation to DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018), demonstrating the generality of our observations. As shown in Figure 2, deeper networks trained with PQN exhibit a collapse in both episode returns and gradient norms [^1] Unless otherwise specified, all ALE results are averaged over three seeds., highlighting the fragility of deep models under non-stationarity.
@@ 3.2 Training Degradation
In Figure 3 we evaluate diagnostic metrics capturing expressivity and training dynamics, revealing that deeper networks exhibit pronounced training pathologies and degraded performance. We first measure the fraction of dormant neurons, defined as units with near-zero activations over a batch of trajectories (Sokar et al., 2023), and find that dormant neurons grow with depth, signaling underutilized capacity. Next, we assess representational expressivity using SRank, the effective rank of penultimate-layer activations (Kumar et al., 2020), observing that deeper networks tend to collapse state representations into lower-dimensional, and less expressive (as evidenced by declining returns) subspaces. To study loss curvature, we compute the Hessian trace of the temporal-difference loss. This metric serves as a proxy for sharpness or smoothness in optimization (Ghorbani et al., 2019), similarly to tracking the largest eigenvalue. Figure 3 shows that only shallow networks exhibit high Hessian trace values, suggesting access to sharper regions of the loss surface with pronounced directions of improvement. In contrast, deeper architectures consistently show near-zero trace, indicating poorly conditioned geometry that hinders effective gradient-based updates. These findings suggest a breakdown in representation, plasticity, and optimization as networks scale, ultimately impeding learning.
@@ 4 Stabilizing Gradients
Having identified the pathologies that emerge in non-stationary regimes, particularly under large-scale architectures, we investigate strategies to mitigate these instabilities. We focus on two complementary interventions: skip connections (He et al., 2016) and optimizers (Martens and Grosse, 2015), as these directly improve gradient flow. We continue to use PQN as our base RL algorithm and evaluate on the Atari-10 suite (Aitchison et al., 2023). In section 5, we demonstrate that the effectiveness of our proposed gradient interventions generalize beyond this specific algorithm and environment suite.
@@ 4.1 Intervention 1: Multi-Skip Residuals for Gradient Stability
We compare our network architecture against the standard fully connected baseline across varying depths. As shown in Figure 5 (left), performance collapses with increased depth in the baseline, while the multi-skip architecture maintains stable learning and continues to improve across widths. This improvement is accompanied by consistently higher gradient magnitudes. Complete results across all network depths and widths are presented in subsection C.3.
@@ 4.2 Intervention 2: Second-Order Optimizers for Non-Stationarity
Kronecker-factored optimizer (Kron for short) approximates the FIM and applies structured preconditioning that captures inter-parameter dependencies, unlike Adam’s diagonal scaling. This yields directionally aware preconditioning that better aligns with the curvature of the loss surface (Martens, 2020). In non-stationary settings, such as deep RL, where both the data distribution and curvature evolve over time, curvature-aware updates can help preserve gradient signal by maintaining stable update magnitudes and directions. As shown in Figure 5 (right), replacing RAdam with Kron prevents performance collapse at greater depths, even in standard MLP architectures. Complete results across all network depths and widths are presented in subsection C.3.
@@ 4.3 Combining Gradient Interventions
We combine both gradient interventions to PQN and evaluate it on the full ALE suite (57 games), across 3 seeds and 200M frames. Figure 6 shows that our augmented agent outperforms the baseline in 90% of the environments, achieving a median relative improvement of 83.27%. Notably, the baseline PQN is itself competitive with strong agents such as Rainbow (Gallici et al., 2025), highlighting the effectiveness of our interventions. Detailed per-environment learning curves can be found in subsection C.4.
In Figure 7 we validate the effectiveness of the combined gradient interventions in the non-stationary SL setting we used as motivation in section 3. The results verify that these interventions enable high accuracy and sustained adaptability across depths and widths, even under dynamic label reshuffling.
@@
Figure 6: Gradient-stabilized PQN achieves superior scalability. (Left) On Atari-10, the combined interventions lead to high HNS even at greater depths, outperforming either intervention alone (see Figure 5) and increased gradient gradient flow. (Right) On the full ALE suite, our agent outperforms the baseline in 90% of the games with a median performance improvement of 83.27%.

### v2

@@ Abstract
Scaling deep reinforcement learning networks is challenging and often results in degraded performance, yet the root causes of this failure mode remain poorly understood. Several recent works have proposed mechanisms to address this, but they are often complex and fail to highlight the causes underlying this difficulty. In this work, we conduct a series of empirical analyses which suggest that the combination of non-stationarity with gradient pathologies, due to suboptimal architectural choices, underlie the challenges of scale. We propose a series of direct interventions that stabilize gradient flow, enabling robust performance across a range of network depths and widths. Our interventions are simple to implement and compatible with well-established algorithms, and result in an effective mechanism that enables strong performance even at large scales. We validate our findings on a variety of agents and suites of environments. Source code here.
@@ 1 Introduction
Recent advances in deep reinforcement learning (deep RL) have demonstrated the ability of deep neural networks to solve complex decision-making tasks from robotics to game play and resource optimization (Mnih et al., 2015; Vinyals et al., 2019; Bellemare et al., 2020; Fawzi et al., 2022). Motivated by successes in supervised and generative learning, recent works have explored scaling architectures in deep RL, showing gains in representation quality and generalization across tasks (Farebrother et al., 2023; Taiga et al., 2023). However, scaling neural networks in deep RL remains fundamentally challenging (Ceron et al., 2024c, b). A central cause of this instability lies in the unique optimization challenges of RL. Unlike supervised learning, where data distributions are fixed, deep RL involves policy-dependent data that constantly change during training (Lyle et al., 2022). Each update of the policy $\pi_{\theta}$ alters future states and rewards, making the training objective inherently non-stationary. Value-based methods exacerbate these issues via bootstrapping, recursively using predicted values as targets.
Estimation errors compound over time (Fujimoto et al., 2018), especially under sparse or delayed rewards (Zheng et al., 2018), leading to unstable updates, policy collapse, or value divergence (Van Hasselt et al., 2016; Lyle et al., 2023, 2024). These dynamics are tightly coupled with architectural vulnerabilities. Deep networks face well known pathologies such as vanishing/exploding gradients (Pascanu et al., 2013), ill-conditioned Jacobians (Pennington et al., 2017), and activation saturation (Glorot and Bengio, 2010). In deep RL, these are magnified by the “deadly triad(Sutton and Barto, 2018; Van Hasselt et al., 2018), off-policy corrections, and changing targets. As networks scale, the risk of signal distortion and misalignment increases, resulting in underutilized capacity and brittle learning (Obando Ceron et al., 2023; Ceron et al., 2024b).
One overlooked source of these failures lies in how gradients propagate through the network. Specifically, the gradient decomposition, the layer-wise structure of backpropagation as a chain of Jacobians and weights determine how information flows during learning (Lee et al., 2020). While gradient signal preservation has been studied in supervised learning (Schoenholz et al., 2017; Jacot et al., 2018), its role in deep RL, where both inputs and targets shift continually, remains poorly understood.
@@ 2 Preliminaries
A deep reinforcement learning agent interacts with an environment through sequences of actions ($a\in\mathcal{A}$), which produce corresponding sequences of observations ($s\in\mathcal{S}$) and rewards ($r\in\mathbb{R}$), resulting in trajectories of the form $\tau:=\{s_{0},a_{0},r_{0},s_{1},a_{1},r_{2},\ldots\}$. The agent’s behavior is often represented by a neural network with parameters $\theta$, composed of convolutional layers $\{\phi_{1},\phi_{2},\dots,\phi_{L_{c}}\}$ and dense (fully connected) layers $\{\psi_{1},\psi_{2},\dots,\psi_{L_{d}}\}$, where $\psi_{L_{d}}$ has an output dimensionality of $|\mathcal{A}|$. At every timestep $t$, an observation $s_{t}\in\mathcal{S}$ is fed through the network to obtain an estimate of the long-term value of each action: $Q_{\theta}(s_{t},\cdot)=\psi_{L_{d}}(\psi_{L_{d}-1}(\ldots(\phi_{L_{c}}(\ldots(\phi_{1}(s_{t}))\ldots))\ldots))$. The agent’s policy $\pi_{\theta}(\cdot\mid s_{t})$ specifies the probability of selecting each action, for instance by taking the softmax over the estimated values as in Eq. 1. The training objective is typically defined as the maximization of expected cumulative reward as in Eq. 2,
@@
While training large models in supervised learning settings present challenges, advances in initialization, normalization, and scaling strategies have enabled relatively stable optimization (Ioffe and Szegedy, 2015; Ba et al., 2016; Glorot and Bengio, 2010). Scaling up model size has been a central driver of progress across domains, improving generalization, enhancing representation learning, and boosting downstream performance (Kaplan et al., 2020).
Deep RL differs substantially from supervised learning. First, the data distribution is non-stationary, continually shifting as $\pi_{\theta}$ updates. Second, learning signals are often sparse, delayed, or noisy, which introduces variance in the estimated gradients (Han et al., 2022; Fujimoto et al., 2018; Liu et al., 2025b). These factors destabilize optimization and lead to loss surfaces with sharp curvature and complex local structure (Ilyas et al., 2020; Achiam et al., 2019). Moreover, increasing model capacity often degrades performance unless regularization or architectural interventions are applied (Gogianu et al., 2021; Bjorck et al., 2021; Schwarzer et al., 2023; Wang et al., 2025).
@@
where each $\frac{\partial h_{k}}{\partial h_{k-1}}\in\mathbb{R}^{d_{k}\times d_{k-1}}$ is the Jacobian. If the singular values of these Jacobians are not properly controlled, their repeated multiplication can cause the norm of the gradient to shrink or grow exponentially with $\mathcal{L}$. This severely impairs convergence, as earlier layers receive little to no useful gradient signal or become numerically unstable (Ioffe and Szegedy, 2015; He et al., 2016).
@@
where $\zeta^{\prime}(\cdot)$ denotes the elementwise derivative of the activation function, and $\odot$ represents elementwise multiplication. For ReLU, $\zeta^{\prime}(x)=\mathbf{1}_{x>0}$, so the gradient is entirely blocked wherever the neuron is inactive. This leads to the well-known dying ReLU problem, where a significant portion of the network ceases to update and becomes untrainable (Lu et al., 2019; Shin and Karniadakis, 2020).
@@ 3 Diagnosis: Gradients Under Non-stationarity and Scale
A fundamental premise of modern deep learning is that scaling model capacity yields consistent gains in performance (Kaplan et al., 2020; Chowdhery et al., 2023). This has held true in large-scale supervised learning, where training data distributions are stationary and i.i.d., and gradient descent operates under relatively stable conditions. However, in non-stationary settings, such as RL, gradient-based optimization faces severe challenges that scaling alone may exacerbate (Ceron et al., 2024b, c). In this section, we diagnose how gradient pathologies emerge and intensify across different settings, with a focus on architectural scaling in width and depth (network scales used specified in Table 1).
@@ 3.1 Gradient Pathologies
Supervised Learning (Stationary and Non-Stationary) We use the CIFAR-10 image classification benchmark (Krizhevsky and others, 2009), where the input-output mapping remains fixed over time. Models consist of standard 6-layer convolutional neural networks (CNN) followed by a multi-layer perceptron (MLP). We vary the depth and width of the MLP to explore how model scale influences learning behavior. To introduce non-stationarity, we periodically shuffle the training labels during training, following the setup by Sokar et al. (2023). This creates a loss landscape that changes over time, echoing the challenges of deep RL. Fig. 1 illustrates the contrast in training behavior and gradient flow between stationary and non-stationary supervised learning. Under non-stationarity, deep networks fail to recover accuracy, which aligns with a marked degradation in gradient magnitudes.
@@
Reinforcement Learning As discussed in Sec. 2, RL introduces fundamentally different sources of non-stationarity due to the policy-dependent data distribution and moving target estimates. To study gradient dynamics, we use PQN (Gallici et al., 2025), a recent value-based algorithm that achieves strong performance without relying on a target network or replay buffer. PQN ensures stability and convergence using Layer Normalization (Ba et al., 2016) and supports GPU-based training through vectorized environments for online parallel data collection. In Sec. C.1 we extend our investigation to DQN (Mnih et al., 2015) and Rainbow (Hessel et al., 2018), demonstrating the generality of our observations. As shown in Fig. 2, deeper networks trained with PQN exhibit a collapse in both episode returns and gradient norms [^1] Unless otherwise specified, all ALE results are averaged over three seeds., highlighting the fragility of deep models under non-stationarity.
@@ 3.2 Training Degradation
In Fig. 3 we evaluate diagnostic metrics capturing expressivity and training dynamics, revealing that deeper networks exhibit pronounced training pathologies and degraded performance. We first measure the fraction of dormant neurons, defined as units with near-zero activations over a batch of trajectories (Sokar et al., 2023), and find that dormant neurons grow with depth, signaling underutilized capacity. Next, we assess representational expressivity using SRank, the effective rank of penultimate-layer activations (Kumar et al., 2020), observing that deeper networks tend to collapse state representations into lower-dimensional, and less expressive (as evidenced by declining returns) subspaces. To study loss curvature, we compute the Hessian trace of the temporal-difference loss. This metric serves as a proxy for sharpness or smoothness in optimization (Ghorbani et al., 2019), similarly to tracking the largest eigenvalue. Fig. 3 shows that only shallow networks exhibit high Hessian trace values, suggesting access to sharper regions of the loss surface with pronounced directions of improvement. In contrast, deeper architectures consistently show near-zero trace, indicating poorly conditioned geometry that hinders effective gradient-based updates. These findings suggest a breakdown in representation, plasticity, and optimization as networks scale, ultimately impeding learning.
@@ 4 Stabilizing Gradients
Having identified the pathologies that emerge in non-stationary regimes, particularly under large-scale architectures, we investigate strategies to mitigate these instabilities. We focus on two complementary interventions: skip connections (He et al., 2016) and optimizers (Martens and Grosse, 2015), as these directly improve gradient flow. We continue to use PQN as our base RL algorithm and evaluate on the Atari-10 suite (Aitchison et al., 2023). In Sec. 5, we demonstrate that the effectiveness of our proposed gradient interventions generalize beyond this specific algorithm and environment suite.
@@ 4.1 Intervention 1: Multi-Skip Residuals for Gradient Stability
We compare our network architecture against the standard fully connected baseline across varying depths. As shown in Fig. 5 (left), performance collapses with increased depth in the baseline, while the multi-skip architecture maintains stable learning and continues to improve across widths. This improvement is accompanied by consistently higher gradient magnitudes. Complete results across all network depths and widths are presented in Sec. C.4.
@@ 4.2 Intervention 2: Second-Order Optimizers for Non-Stationarity
Kronecker-factored optimizer (Kron for short) approximates the FIM and applies structured preconditioning that captures inter-parameter dependencies, unlike Adam’s diagonal scaling. This yields directionally aware preconditioning that better aligns with the curvature of the loss surface (Martens, 2020). In non-stationary settings, such as deep RL, where both the data distribution and curvature evolve over time, curvature-aware updates can help preserve gradient signal by maintaining stable update magnitudes and directions. As shown in Fig. 5 (right), replacing RAdam with Kron prevents performance collapse at greater depths, even in standard MLP architectures. Complete results across all network depths and widths are presented in Sec. C.6.
@@ 4.3 Combining Gradient Interventions
We combine both gradient interventions to PQN and evaluate it on the full ALE suite (57 games), across 3 seeds and 200M frames. Fig. 6 shows that our augmented agent outperforms the baseline in 90% of the environments, achieving a median relative improvement of 83.27%. Notably, the baseline PQN is itself competitive with strong agents such as Rainbow (Gallici et al., 2025), highlighting the effectiveness of our interventions. Detailed per-environment learning curves can be found in Sec. E.1.
In Fig. 7 we validate the effectiveness of the combined gradient interventions in the non-stationary SL setting we used as motivation in Sec. 3. The results verify that these interventions enable high accuracy and sustained adaptability across depths and widths, even under dynamic label reshuffling.
@@
Figure 6: Gradient-stabilized PQN achieves superior scalability. (Left) On Atari-10, the combined interventions lead to high HNS even at greater depths, outperforming either intervention alone (see Fig. 5) and increased gradient gradient flow. (Right) On the full ALE suite, our agent outperforms the baseline in 90% of the games with a median performance improvement of 83.27%.
@@
4.4 Alternative Gradient-Stabilization Methods
To ensure that our findings were not specific to a narrow choice of interventions, we conducted a broader exploration of alternative strategies for improving gradient stability in deep RL. We tested a variety of approaches inspired by prior work on optimization and representation stability in both supervised and RL settings (see Sec. C.7 for more details).
As summarized in Tab. 3, none of these methods consistently improved stability or performance compared to our proposed combination of multi-skip residuals and Kronecker-factored optimization. In many cases, the alternatives yielded either negligible gains or degraded performance as network depth increased, reinforcing that architectural and curvature-aware interventions are key to preserving gradient flow at scale.